# TCR-EML: Explainable Model Layers for TCR-pMHC Prediction

### v1

@@ 1 Introduction
For the adaptive immune system, T cells are essential for detecting and responding to antigens from pathogens such as viruses, bacteria, and cancer cells (Joglekar & Li, 2021), as well as in autoimmune contexts. The final step of T cell activation involves binding between a peptide presented by the Major Histocompatibility Complex (pMHC) and the T cell receptor (TCR). The specificity of this interaction is the foundation of T cell-mediated immunity and is a major focus of research in both therapeutic development and the study of immune mechanisms. A detailed understanding of T cell response is critical for designing vaccines that provide durable immunity and for developing effective personalized cancer treatments (Rojas et al., 2023; Poorebrahim et al., 2021).
CD8+ T cells are activated through the MHCI pathway, whereas CD4+ T cells are activated through the MHCII pathway. Epitope prediction for CD8+ T cells has achieved notable success, while mechanisms of CD4+ T cell response remain less well understood. The CD4+ T cell response can be viewed as a two-stage recognition process. In the first stage, antigens are processed by antigen-presenting cells (APCs) and loaded onto MHCII molecules, which are subsequently presented on the APC surface (Davis & Bjorkman, 1988; Neefjes et al., 2011). In the second stage, TCRs recognize these pMHC complexes, initiating the T cell response. TCR recognition is mediated by the $\alpha$ and $\beta$ domains, each composed of variable (V), joining (J), and constant (C) regions, with the $\beta$ chain also containing a diversity (D) region (Bosselut, 2019). Accurate prediction of T cell responses therefore requires modeling both antigen processing and TCR-pMHC binding (Peters et al., 2020; Nielsen et al., 2020).
Early computational work in epitope prediction emphasized peptide-MHCII binding using allele-specific machine learning methods (Nielsen et al., 2020), with tools including NetMHCpan (Hoof et al., 2009; Nielsen et al., 2007) and NetMHCcons (Karosiene et al., 2012). More recently, antigen processing has been modeled with the Antigen Processing Likelihood (APL) algorithm (Mettu et al., 2016; Bhattacharya et al., 2023; Li et al., 2024a; b; Charles et al., 2022), which accounts for structural features that determine which peptides are presented by MHCII molecules.
@@
Prediction of TCR-pMHC binding remains a central challenge in quantitative immunology and adaptive immunity (Hudson et al., 2023). Both unsupervised and supervised learning algorithm approaches have been explored (Hudson et al., 2023; 2024). Unsupervised methods cluster TCR sequences using similarity metrics such as TCRdist3 (Mayer-Blackwell et al., 2021) applied to complementarity-determining regions (CDRs), without requiring binding labels or epitope information (e.g., GIANA (Zhang et al., 2021), GLIPH2 (Huang et al., 2020)). The resulting clusters are then used to support experimental analysis (Hudson et al., 2024). Supervised methods leverage large TCR-pMHC datasets (Hudson et al., 2023) from sources including VDJdb (Bagaev et al., 2020), McPAS-TCR (Tickotsky et al., 2017), and IEDB (Vita et al., 2019), and employ deep learning models such as MixTCRpred (Croce et al., 2024), NetTCR2.2 (Jensen & Nielsen, 2023), TULIP (Meynard-Piganeau et al., 2024), and EGM (Li et al., 2025a).
@@ 4.1 Datasets and Benchmarks
To train and evaluate our model layers, we constructed a TCR-pMHC dataset comprising 349,716 paired samples of TCR alpha and beta chains, 2,316 unique peptides, 29,581 distinct CDR3a sequences, and 32,578 distinct CDR3b sequences. The dataset spans multiple species, primarily Homo sapiens and Mus musculus. Among the samples, 95.7% correspond to MHC-I and 4.3% to MHC-II. The dataset was compiled from VDJdb (Bagaev et al., 2020), TCR-McPAS (Tickotsky et al., 2017), IEDB (Vita et al., 2019), TBAdb Zhang et al. (2020), and 10x Genomics (10x Genomics, 2022). For all sources, we retained only samples that provided CDR3a, CDR3b, and peptide sequences. Any non-standard characters, irregular notations, or missing residues were discarded in the amino acid sequences. Negative samples were generated by directly sampling negative pairs for the 10x Genomics dataset, and for other datasets by randomly shuffling TCR and pMHC pairs. For each epitope, negative samples were generated at a ratio of 4:1 relative to positive samples. Finally, the dataset was split into training and test sets using a 95:5 ratio. The test set contains 15,503 samples spanning 288 epitopes that do not appear in the training data. To construct the evaluation set, we computed the Levenshtein distance between each pair of peptides. We then sampled a comparable number of peptides whose minimal pairwise distance exceeded different thresholds from 1 to 9, ensuring that all selected epitopes were excluded from and distinct from those in the training dataset.
@@ 4.2 Protein Language Models and Baselines
We consider two categories of comparable models in our evaluations. First, to evaluate PLM-based models, we constructed a standard linear classifier for PLM features. Specifically, we added two fully connected layers with hidden dimension three times the feature dimension and ReLU activation. Each classifier takes concatenated global representations of CDR3a, CDR3b, and the peptide as input and outputs a prediction score. For ProteinBERT, we used the provided global features, and for the ESM models, we averaged local residue-level features to obtain global representations. Second, we compared our models with two recent transformer-based TCR-pMHC prediction methods. MixTCRpred (Croce et al., 2024), one of the most widely used TCR-pMHC models, utilizes all CDR regions as input. We evaluated MixTCRpred on only CDR3 regions. TULIP (Meynard-Piganeau et al., 2024) is another recent model that outperforms the widely used NetTCR-2.2 (Jensen & Nielsen, 2023) baseline in terms of accuracy and generalization ability.

### v2

@@ 1 Introduction
For the adaptive immune system, T cells are essential for detecting and responding to antigens from pathogens such as viruses, bacteria, and cancer cells (Joglekar and Li, 2021), as well as in autoimmune contexts. The final step of T cell activation involves binding between a peptide presented by the Major Histocompatibility Complex (pMHC) and the T cell receptor (TCR). The specificity of this interaction is the foundation of T cell-mediated immunity and is a major focus of research in both therapeutic development and the study of immune mechanisms. A detailed understanding of T cell response is critical for designing vaccines that provide durable immunity and for developing effective personalized cancer treatments (Rojas et al., 2023; Poorebrahim et al., 2021).
CD8+ T cells are activated through the MHCI pathway, whereas CD4+ T cells are activated through the MHCII pathway. Epitope prediction for CD8+ T cells has achieved notable success, while mechanisms of CD4+ T cell response remain less well understood. The CD4+ T cell response can be viewed as a two-stage recognition process. In the first stage, antigens are processed by antigen-presenting cells (APCs) and loaded onto MHCII molecules, which are subsequently presented on the APC surface (Davis and Bjorkman, 1988; Neefjes et al., 2011). In the second stage, TCRs recognize these pMHC complexes, initiating the T cell response. TCR recognition is mediated by the $\alpha$ and $\beta$ domains, each composed of variable (V), joining (J), and constant (C) regions, with the $\beta$ chain also containing a diversity (D) region (Bosselut, 2019). Accurate prediction of T cell responses therefore requires modeling both antigen processing and TCR-pMHC binding (Peters et al., 2020; Nielsen et al., 2020).
Early computational work in epitope prediction emphasized peptide-MHCII binding using allele-specific machine learning methods (Nielsen et al., 2020), with tools including NetMHCpan (Hoof et al., 2009; Nielsen et al., 2007) and NetMHCcons (Karosiene et al., 2012). More recently, antigen processing has been modeled with the Antigen Processing Likelihood (APL) algorithm (Mettu et al., 2016; Bhattacharya et al., 2023; Li et al., 2024b; a; Charles et al., 2022), which accounts for structural features that determine which peptides are presented by MHCII molecules.
@@
Prediction of TCR-pMHC binding remains a central challenge in quantitative immunology and adaptive immunity (Hudson et al., 2023). Both unsupervised and supervised learning algorithm approaches have been explored (Hudson et al., 2023; 2024). Unsupervised methods cluster TCR sequences using similarity metrics such as TCRdist3 (Mayer-Blackwell et al., 2021) applied to complementarity-determining regions (CDRs), without requiring binding labels or epitope information (e.g., GIANA (Zhang et al., 2021), GLIPH2 (Huang et al., 2020)). The resulting clusters are then used to support experimental analysis (Hudson et al., 2024). Supervised methods leverage large TCR-pMHC datasets (Hudson et al., 2023) from sources including VDJdb (Bagaev et al., 2020), McPAS-TCR (Tickotsky et al., 2017), and IEDB (Vita et al., 2019), and employ deep learning models such as MixTCRpred (Croce et al., 2024), NetTCR2.2 (Jensen and Nielsen, 2023), TULIP (Meynard-Piganeau et al., 2024), and EGM (Li et al., 2025a).
@@ 4.1 Datasets and Benchmarks
To train and evaluate our model layers, we constructed a TCR-pMHC dataset comprising 349,716 paired samples of TCR alpha and beta chains, 2,316 unique peptides, 29,581 distinct CDR3a sequences, and 32,578 distinct CDR3b sequences. The dataset spans multiple species, primarily Homo sapiens and Mus musculus. Among the samples, 95.7% correspond to MHC-I and 4.3% to MHC-II. The dataset was compiled from VDJdb (Bagaev et al., 2020), TCR-McPAS (Tickotsky et al., 2017), IEDB (Vita et al., 2019), TBAdb Zhang et al. (2020), and 10x Genomics (Genomics, 2022). For all sources, we retained only samples that provided CDR3a, CDR3b, and peptide sequences. Any non-standard characters, irregular notations, or missing residues were discarded in the amino acid sequences. Negative samples were generated by directly sampling negative pairs for the 10x Genomics dataset, and for other datasets by randomly shuffling TCR and pMHC pairs. For each epitope, negative samples were generated at a ratio of 4:1 relative to positive samples. Finally, the dataset was split into training and test sets using a 95:5 ratio. The test set contains 15,503 samples spanning 288 epitopes that do not appear in the training data. To construct the evaluation set, we computed the Levenshtein distance between each pair of peptides. We then sampled a comparable number of peptides whose minimal pairwise distance exceeded different thresholds from 1 to 9, ensuring that all selected epitopes were excluded from and distinct from those in the training dataset.
@@ 4.2 Protein Language Models and Baselines
We consider two categories of comparable models in our evaluations. First, to evaluate PLM-based models, we constructed a standard linear classifier for PLM features. Specifically, we added two fully connected layers with hidden dimension three times the feature dimension and ReLU activation. Each classifier takes concatenated global representations of CDR3a, CDR3b, and the peptide as input and outputs a prediction score. For ProteinBERT, we used the provided global features, and for the ESM models, we averaged local residue-level features to obtain global representations. Second, we compared our models with two recent transformer-based TCR-pMHC prediction methods. MixTCRpred (Croce et al., 2024), one of the most widely used TCR-pMHC models, utilizes all CDR regions as input. We evaluated MixTCRpred on only CDR3 regions. TULIP (Meynard-Piganeau et al., 2024) is another recent model that outperforms the widely used NetTCR-2.2 (Jensen and Nielsen, 2023) baseline in terms of accuracy and generalization ability.
@@ 5 Conclusion
Acknowledgments
The authors acknowledge support from the Harold L. and Heather E. Jurist Center of Excellence for Artificial Intelligence at Tulane University. This work was also supported by the National Institutes of Health (U54-CA260581) through the Tulane University COVID Antibody and Immunity Network (TUCAIN); Tulane SOM Pilot Funding for “MHCII Pathway Processing of SARS-CoV-2 Spike”; the Tulane Center of Excellence for Emerging and Re-emerging Infectious Diseases (CEERID) Pilot Research Program for “Large-Scale Validation of a Novel Parallel Algorithm for Computational Epitope Prediction”; and the Lavin-Bernick Faculty Grant Proposal Research and Scholarly Activities Support for “Research Trainee Support for Modeling Antigen Processing and HLA Immunopeptidomics”.